USING A COMBINATION OF 24% NITROGEN AND 12% NITROGEN TO ACHIEVE AN END RESULT OF 100 GRAMS OF 21% NITROGEN.
Answer by Earlsdon(6294)

Set up your problem this way:
Assign the variable(s):
Let x = # of grams of 24% nitrogen and 100 - x = # of grams of 12% nitrogen.
Convert the percents to decimals.
x(.24)+ (100 - x)(.12) = 100((.21) Simplify.
.24x + 12 - .12x = 21
.12x = 9
x = 9/(.12)
x = 75 grams
100 - x = 100 - 75 = 25 grams
You would need 75 grams of 24% nitrogen plus 25 grams of 12% nitrogen to achieve a mixture of 100 grams of 21% nitrogen.